# The First and Second-Order Asymptotics of Covert Communication over AWGN Channels

### v4

@@ IV-B Resolvability Analysis
[…]
$\mathbb{E}$ $\left[\int_{\bm{z}}\hat{Q}^{(n)}_{\bm{Z}}(\bm{z})\log\frac{\hat{Q}^{(n)}_{1}(\bm{z})}{\bar{Q}^{(n)}_{1}(\bm{z})}d\bm{z}\right]$
$=$ $\int_{\bm{z}}\left(\int_{\bm{x}}\frac{1}{MK}\sum_{s=1}^{K}\sum_{w=1}^{M}W^{n}_{2}(\bm{z}|\bm{x}_{k,i})d\bar{P}_{n}^{(n)}\right)$
$\times$ $\log\frac{\hat{Q}^{(n)}_{1}(\bm{z})}{\bar{Q}^{(n)}_{1}(\bm{z})}d\bm{z}$
=$\overset{(a)}{=}$ $\int_{\bm{z}}\hat{Q}^{(n)}_{1}(\bm{z})\log\frac{\hat{Q}^{(n)}_{1}(\bm{z})}{\bar{Q}^{(n)}_{1}(\bm{z})}d\bm{z}$
$=$ $\mathbb{D}(\hat{Q}^{(n)}_{1}\|\bar{Q}^{(n)}_{1}).$ (82)
In step (a), we have used the fact that each $\bm{x}_{k,i}$ are i.i.d generated from $\bar{P}_{n}^{(n)}$ and the fact that the induced distribution of $\bar{P}_{n}^{(n)}$ is $\hat{Q}^{(n)}_{1}$. For the third term in (63), since the distributions $\bar{Q}^{(n)}_{1}$ and $\bar{Q}^{(n)}_{0}$ are independent of $\bm{x}$, we have
$\mathbb{E}$ $\left[\int_{\bm{z}}\hat{Q}^{(n)}_{\bm{Z}}(\bm{z})\log\frac{\bar{Q}^{(n)}_{1}(\bm{z})}{\bar{Q}^{(n)}_{0}(\bm{z})}d\bm{z}\right]$
$=$ $\int_{\bm{z}}\left(\int_{\bm{x}}\frac{1}{MK}\sum_{s=1}^{K}\sum_{w=1}^{M}W^{n}_{2}(\bm{z}|\bm{x}_{k,i})d\bar{P}_{n}^{(n)}\right)$
$\times$ $\log\frac{\bar{Q}^{(n)}_{1}(\bm{z})}{\bar{Q}^{(n)}_{0}(\bm{z})}d\bm{z}$
=$\overset{(a)}{=}$ $\int_{\bm{z}}\hat{Q}^{(n)}_{1}(\bm{z})\log\frac{\bar{Q}^{(n)}_{1}(\bm{z})}{\bar{Q}^{(n)}_{0}(\bm{z})}d\bm{z}$
$=$ $\int_{\bm{z}}\frac{\hat{Q}^{(n)}_{1}(\bm{z})}{\bar{Q}^{(n)}_{1}(\bm{z})}\cdot\bar{Q}^{(n)}_{1}(\bm{z})\log\frac{\bar{Q}^{(n)}_{1}(\bm{z})}{\bar{Q}^{(n)}_{0}(\bm{z})}d\bm{z}$
$\overset{(b)}{\leq}$ $(1+\iota^{\prime})\int_{\bm{z}}\bar{Q}^{(n)}_{1}(\bm{z})\log\frac{\bar{Q}^{(n)}_{1}(\bm{z})}{\bar{Q}^{(n)}_{0}(\bm{z})}d\bm{z}$
$=$ $(1+\iota^{\prime})\mathbb{D}(\bar{Q}^{(n)}_{1}\|\bar{Q}^{(n)}_{0}).$ (83)
In the above derivation, the step (a) follows from $\hat{Q}^{(n)}_{1}$ is the induced distribution of $\bar{P}_{n}^{n}$ through the channel between Alice and Willie, step (b) follows from Lemma 3.
From the above (IV-B), (IV-B) and (IV-B), we have
$\mathbb{E}$ $\left[\mathbb{D}(\hat{Q}^{(n)}_{\bm{z}}||\hat{Q}^{n}_{0})\right]$
@@
$+$ $\mathbb{D}(\hat{Q}^{(n)}_{1}\|\bar{Q}^{(n)}_{1})+(1+\iota^{\prime})\mathbb{D}(\bar{Q}^{(n)}_{1}\|\bar{Q}^{(n)}_{0})$
@@
$+$ $\mathbb{D}(\bar{P}_{n}^{n}\|P^{n}_{n})+(1+\iota^{\prime})\mathbb{D}(\bar{Q}^{(n)}_{1}\|\bar{Q}^{(n)}_{0})$ (84)
where step (a) follows from data processing inequality and the fact that $\hat{Q}^{(n)}_{1}$ and $\bar{Q}^{(n)}_{1}$ are the induced output distributions of $\bar{P}_{n}^{n}$ and $P^{n}_{n}$ through the channel $W_{2}^{n}$, respectively. ∎
@@
In the resolvability analysis, we directly analyze $\mathbb{D}(\hat{Q}^{(n)}_{\bm{z}}||\hat{Q}^{n}_{0})$. This is in contrast to the case in DMCs in [23] where the expectation of the KL divergence between the mixture distribution and the induced distribution of the generating distribution is bounded. This is because that the modified triangle inequality does not hold for KL divergence for distributions in $\mathbb{R}^{n}$.
@@
In the upper bound of (81), the first term represents effect of randomly sampling finite codewords. We can see that when there are more codewords, the ratio of it in the upper bound will be smaller. Moreover, the power level of the codewords and the noise at the adversary are also involved with the first term, we will explain it in detail later. The second term represents the effect of truncation in the generating distribution with respect to isotropic Gaussian distributions, which is the cost for necessary power constraint of finite blocklength random coding. The third term represents KL divergence between two isotropic Gaussian distributions, in which the power level plays a vital role.
IV-C Existence of the Codebooks and Achievability Bound
@@
In last two sections, we have established the decoding error bounds in (55), (59) and the resolvability bounds in (81). Although the bound on error probability of the first type in (55) is in the form of maximal probability of error, the second type of error probability in (59) and the resolvability bound are the average over the generating distribution $\bar{P}_{n}^{n}$. In this section, we will first combine these results to show that, with positive probability, a new family of codebooks can be obtained by an expurgation operation on the previous codebooks sampled from $\bar{P}_{n}^{n}$, at the mean time, each codebook in this new family has an upper bound on the second type of decoding error $\varepsilon^{(2)}$. Next, we determined various intermediate parameters and provide the size of the codebooks as well as the number of key bits. The roadmap is the same as [23]. Nevertheless, as we are dealing with the AWGN channel, the technical details are different. As we will see, the noise level $\sigma^{2}$ and the hyper-parameter of truncation $\mu$ play a vital role, which are absent from the case of DMCs.
@@
The following lemma is necessary for the application of McDiarmid’s Inequality for the codewords over the AWGN channel.
@@
Let $\mathcal{C}_{1}\triangleq\left\{\bm{x}_{1},\cdots,\bm{x}_{k},\cdots,\bm{x}_{M}\right\}$ and $\mathcal{C}_{2}\triangleq\{\bm{x}_{1},\cdots,\bm{x}^{\prime}_{k},\cdots,\bm{x}_{M}\}$ be two codebooks with only one difference on $k$-th codeword. The codewords are of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are all in the support set of $\bar{P}_{n}^{n}$. If $Q_{\bm{Z}}^{1}$ and $Q_{\bm{Z}}^{2}$ are the mixture Gaussian distributions induced by $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ through the channel $W_{2}$, respectively. The following bound holds for the difference between $\mathbb{D}(Q_{\bm{Z}}^{1}\|\bar{Q}^{(n)}_{0})$ and $\mathbb{D}(Q_{\bm{Z}}^{2}\|\bar{Q}^{(n)}_{0})$.
$\begin{split}\big{|}\mathbb{D}(Q_{\bm{Z}}^{1}\|\bar{Q}^{(n)}_{0})&-\mathbb{D}(Q_{\bm{Z}}^{2}\|\bar{Q}^{(n)}_{0})\big{|}\leq\frac{4n\Psi(n)}{M\sigma^{2}}.\end{split}$ (85)
@@
Consider the AWGN channel $W_{2}$ defined in (6), if $\{\bm{x}_{w},w=1,\cdots,M\}$ are $M$ codewords randomly sampled from $\bar{P}_{n}^{n}$ and $Q^{(n)}_{\bm{Z}}(\bm{Z})=\frac{1}{M}\sum_{w=1}^{M}W_{2}(\bm{Z}|\bm{x}_{w})$ is the induced mixture distribution of the codebook, then
$\mathbb{P}\bigg{(}\mathbb{D}(Q^{(n)}_{\bm{Z}}\|\bar{Q}^{(n)}_{0})-\mathbb{E}\big{(}\mathbb{D}(Q^{(n)}_{\bm{Z}}\|\bar{Q}^{(n)}_{0})\big{)}\geq\lambda\bigg{)}$
@@
The following lemma is almost a restatement of Lemma 4 in [23]. It provide a sufficient condition that we can expurgate part of the codewords so that the remaining codewords constitute a family of codebooks such that its induced distribution is close to the distribution of noise, meanwhile, each of the codebook has low decoding error probability of the second type.
@@
With AWGN channels $W_{1}$ and $W_{2}$ defined in (5) and (6), respectively, let the codewords $\bm{x}_{sw}$ be i.i.d sampled from $\bar{P}_{n}^{n}$, and denote $\epsilon^{(1)}_{sw}$ and $\epsilon^{(2)}_{sw}$ as the decoding error probability of the first type and second type as (31) and (32). For any subset $\mathcal{I}\subset\langle M\rangle\times\langle K\rangle$, define a mixture distribution $Q_{\bm{Z}}^{\mathcal{I}}$ as
$Q_{\bm{Z}}^{\mathcal{I}}\triangleq\frac{1}{|\mathcal{I}|}\sum_{(s,w)\in\mathcal{I}}W^{n}_{2}(\bm{Z}|\bm{x}_{sw})$ (87)
For this mixture distribution, define events
[…]
@@
$\begin{split}\mathcal{E}_{2}\triangleq&\left\{\forall\mathcal{I}\subset\langle K\rangle\times\langle M\rangle\,\ \,\text{with}\,\ \,|\mathcal{I}|=\lambda_{2}MK,\right.\\ \mathbb{D}(Q_{\bm{Z}}^{\mathcal{I}}\|\bar{Q}^{(n)}_{0})&\leq\log\big{(}1+\frac{2^{\gamma_{1}+1}}{\lambda_{2}MK}\big{)}+An\cdot ae^{-bn}\\ +&\left.\mathbb{D}(\bar{P}_{n}^{n}\|P^{n}_{n})+2\mathbb{D}(\bar{Q}^{(n)}_{1}\|\bar{Q}^{(n)}_{0})+\lambda_{3}\right\}.\end{split}$ (89)
@@
[…]
$=$ $\bigg{(}1-\mathbb{P}\big{(}\frac{1}{M}\sum_{w=1}^{M}\varepsilon_{sw}^{(2)}\geq\lambda_{1}\frac{1}{n^{c_{0}}}\mathbb{E}_{\bar{Q}^{(n)}_{m,1}}\big{(}\frac{\bar{Q}^{(n)}_{m,1}}{Q^{(n)}_{m,1}}\big{)}\big{)}\bigg{)}^{K}$
@@
$\begin{split}&\mathbb{P}(\mathcal{E}_{1}\cap\mathcal{E}_{2})\\ =&1-\mathbb{P}(\mathcal{E}^{c}_{1}\cup\mathcal{E}^{c}_{2})\\ =&1-\mathbb{P}(\mathcal{E}^{c}_{1})-\mathbb{P}(\mathcal{E}^{c}_{2})\\ =&\mathbb{P}(\mathcal{E}_{1})+\mathbb{P}(\mathcal{E}_{1})-1\\ \geq&(1-\frac{1}{\lambda_{1}})^{K}-\exp(MK\mathbb{H}_{b}(\lambda_{2}))\times\exp\bigg{(}-\frac{\lambda_{2}\lambda_{3}^{2}MK\sigma^{4}}{8n^{2}\Psi^{2}(n)}\bigg{)}\end{split}$ (93)
@@
Note that as we have derived upper bound on the maximal decoding error probability of the first type in Lemma 5, the event $\mathcal{E}_{1}$ bounds the maximal average error probability of the second type for all sub-codebooks. From Lemma 12, it is possible to expurgate part of the codewords so that the remaining codewords form a new family of codebooks. Each member of this new family is a reliable code while the induced distribution of the whole family satisfies the covert constraint.
The next theorem present our main results. We will provide characterizations on the size $M_{n}$ of each codebook for covert communication over AWGN channels with sufficient large $n$ through proper choice of the power level $\Psi(n)$. The theorem also characterizes the number of key bits $K_{n}$ that are necessary, and it is shown that $K_{n}$ depends not only on the noise level $\sigma^{2}$, but also the hyper-parameter of truncation $\mu$.
@@
For covert communication over AWGN channels $W_{1}$ and $W_{2}$, with sufficiently large $n$, a sufficient condition for the following three requirements
@@
    Reliability
    $\epsilon_{1}+\frac{\lambda_{1}}{1-\lambda_{2}}\frac{1}{n^{c_{0}}}\mathbb{E}_{\bar{Q}^{(n)}_{m,1}}\big{(}\frac{\bar{Q}^{(n)}_{m,1}}{Q^{(n)}_{m,1}}\big{)}\leq\epsilon,$ (95)
@@
    $\begin{split}\log&\big{(}1+\frac{2^{\gamma_{1}+1}}{\lambda_{2}M_{n}K_{n}}\big{)}+An\cdot ae^{-bn}\\ +&\mathbb{D}(\bar{P}_{n}^{n}\|P^{n}_{n})+(1+\iota^{\prime})\mathbb{D}(\bar{Q}^{(n)}_{1}\|\bar{Q}^{(n)}_{0})+\lambda_{3}\leq\delta.\end{split}$ (96)
@@
    Choose intermediate parameters $\lambda_{1},\lambda_{2},\lambda_{3}$ as (1): $\lambda_{1}=n^{c_{1}}$, (2) $\lambda_{2}=1-n^{-c_{2}}$, (3): $\lambda_{3}=n^{-c_{3}}$ such that $c_{1},c_{2},c_{3}$ all positive, meanwhile, they satisfy $0<c_{1}<c_{0}$, $c_{1}+c_{2}<c_{0}$ and $10<2c_{3}<c_{2}-$ 1.

### v5

@@ III-B Resolvability Analysis
Since we have bounded the first term in Lemma 9, we just consider the second and the third term in (69). For the second term, as the distributions $\bar{Q}^{(n)}_{\bm{Z},1}$ and $Q^{(n)}_{\bm{Z},1}$ are independent of any codeword $\bm{x}$, we have
$\mathbb{E}$ $\left[\int_{\bm{z}}Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})\log\frac{\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})}{{Q}^{(n)}_{\bm{Z},1}(\bm{z})}d\bm{z}\right]$
$=$ $\int_{\bm{z}}\left(\int_{\bm{x}}\frac{1}{MK}\sum_{s=1}^{K}\sum_{w=1}^{M}W^{n}_{2}(\bm{z}|\bm{x}_{sw})d\bar{P}_{n}^{(n)}\right)$
$\times$ $\log\frac{\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})}{{Q}^{(n)}_{\bm{Z},1}(\bm{z})}d\bm{z}$
$\overset{(a)}{=}$ $\int_{\bm{z}}\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})\log\frac{\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})}{{Q}^{(n)}_{\bm{Z},1}(\bm{z})}d\bm{z}$
$=$ $\mathbb{D}(\bar{Q}^{(n)}_{\bm{Z},1}\|{Q}^{(n)}_{\bm{Z},1})$. (91)
In step (a), we have used the fact that each $\bm{x}_{sw}$ are i.i.d generated from $\bar{P}_{n}^{(n)}$ and the fact that the induced distribution of $\bar{P}_{n}^{(n)}$ is $\bar{Q}^{(n)}_{\bm{Z},1}$. For the third term in (69), since the distributions $Q^{(n)}_{\bm{Z},1}$ and $Q^{(n)}_{\bm{Z},0}$ are independent of $\bm{x}$, we have
$\mathbb{E}$ $\left[\int_{\bm{z}}Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})\log\frac{Q^{(n)}_{\bm{Z},1}(\bm{z})}{Q^{(n)}_{\bm{Z},0}(\bm{z})}d\bm{z}\right]$
$=$ $\int_{\bm{z}}\left(\int_{\bm{x}}\frac{1}{MK}\sum_{s=1}^{K}\sum_{w=1}^{M}W^{n}_{2}(\bm{z}|\bm{x}_{sw})d\bar{P}_{n}^{(n)}\right)$
$\times$ $\log\frac{Q^{(n)}_{\bm{Z},1}(\bm{z})}{Q^{(n)}_{\bm{Z},0}(\bm{z})}d\bm{z}$
$\overset{(a)}{=}$ $\int_{\bm{z}}\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})\log\frac{Q^{(n)}_{\bm{Z},1}(\bm{z})}{Q^{(n)}_{\bm{Z},0}(\bm{z})}d\bm{z}$
$=$ $\int_{\bm{z}}\frac{\bar{Q}^{(n)}_{\bm{Z},1}(\bm{z})}{Q^{(n)}_{\bm{Z},1}(\bm{z})}\cdot Q^{(n)}_{\bm{Z},1}(\bm{z})\log\frac{Q^{(n)}_{\bm{Z},1}(\bm{z})}{Q^{(n)}_{\bm{Z},0}(\bm{z})}d\bm{z}$
$\overset{(b)}{\leq}$ $(1+\iota^{\prime})\int_{\bm{z}}Q^{(n)}_{\bm{Z},1}(\bm{z})\log\frac{Q^{(n)}_{\bm{Z},1}(\bm{z})}{Q^{(n)}_{\bm{Z},0}(\bm{z})}d\bm{z}$
$=$ $(1+\iota^{\prime})\mathbb{D}(Q^{(n)}_{\bm{Z},1}(\bm{z})\|Q^{(n)}_{\bm{Z},0}(\bm{z})).$ (92)
In the above derivation, the step (a) follows from $\bar{Q}^{(n)}_{\bm{Z},1}$ is the induced distribution of $\bar{P}_{n}^{n}$ through the channel between Alice and Willie, step (b) follows from Lemma 3.
From the above (III-B), (III-B) and (III-B), we have
$\mathbb{E}$ $\left[\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})\right$]
@@
$+$ $\mathbb{D}(\bar{Q}^{(n)}_{\bm{Z},1}\|{Q}^{(n)}_{\bm{Z},1})+(1+\iota^{\prime})\mathbb{D}(Q^{(n)}_{\bm{Z},1}\|Q^{(n)}_{\bm{Z},0})$
@@
$+$ $\mathbb{D}(\bar{P}_{n}^{(n)}\|P^{(n)}_{n})+(1+\iota^{\prime})\mathbb{D}(Q^{(n)}_{\bm{Z},1}\|Q^{(n)}_{\bm{Z},0})$ (93)
where step (a) follows from data processing inequality and the fact that $\bar{Q}^{(n)}_{\bm{Z},1}$ and $Q^{(n)}_{\bm{Z},1}$ are the induced output distributions of $\bar{P}_{n}^{(n)}$ and $P^{(n)}_{n}$ through the channel $W_{2}^{n}$, respectively. ∎
@@
In the resolvability analysis, we directly analyze the expectation of $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})$ by leveraging intermediate distributions $Q^{(n)}_{\bm{Z},1}$ and $\bar{Q}^{(n)}_{\bm{Z},1}$. The strategy is different from the technique path in DMCs [22], where the divergence between the induced distribution of the code family and the distribution of noise is first bounded using modified triangle inequality, and then each term of the upper bound is separately bounded. This is because that the modified triangle inequality does not hold for KL divergence for distributions in $\mathbb{R}^{n}$.
@@
In the upper bound of (89), the first term represents effect of randomly sampling finite codewords. We can see that when there are more codewords, the ratio of it in the upper bound will be smaller. Moreover, the power level of the codewords and the noise at the adversary are also involved with the first term, we will explain it in detail later. The second term represents the effect of truncation in the generating distribution with respect to isotropic Gaussian distributions, which is the cost for necessary power constraint of finite blocklength random coding. The third term represents KL divergence between two isotropic Gaussian distributions, one of which is the distribution of noise, For the third term, the parameters $\Psi(n)$ and $\sigma$, which represent the power of the codewords and the power of the channel noise of $W_{2}$, are the key elements for the coding scheme and will be determined later.
III-C Existence of the Codebooks and Achievability Bound
@@
In last two sections, we have established the decoding error bounds in (59), (65) and the resolvability bounds in (89). Although the bound on error probability of the first type in (59) is in the form of maximal probability of error, the second type of error probability in (65) and the resolvability bound are the average over the generating distribution $\bar{P}_{n}^{(n)}$. In this section, we will first combine these results to show that, with positive probability, a new family of codebooks can be obtained by an expurgation operation on the previous codebooks sampled from $\bar{P}_{n}^{(n)}$. Each codeword in this new family has an upper bound on the second type of decoding error $\varepsilon^{(2)}$, at the mean time, the induced distribution of the whole family is close to the distribution of the channel noise. As a result, the maximal decoding error probability of each codeword of this new family is low, while the induced distribution of the whole family satisfies the covert constraint. We determined various intermediate parameters and provide the size of the codebooks as well as the number of key bits. The roadmap is the same as [22]. Nevertheless, as we are dealing with the AWGN channel, the technical details are different. As we will see, the noise level $\sigma^{2}$ and the hyper-parameter of truncation $\mu$ play a vital role, which are absent from the case of DMCs.
We first recall McDiarmid’s inequality below.
Theorem 5.
(McDiarmid’s inequality): Let $\{X_{i}\}_{i=1}^{n}$ be independent random variables on some measurable space $\mathcal{X}$, and a measurable function $U=f(X_{1},\cdots,X_{n})$ with $f:\mathcal{X}^{n}\rightarrow\mathbb{R}$ satisfying the following bounded difference assumption:
$\underset{x_{1},\cdots,x_{n},x^{\prime}_{i}\in\mathcal{X}}{\sup}$ $|f(x_{1},\cdots,x_{i},\cdots,x_{n})$
- $f(x_{1},\cdots,x^{\prime}_{i},\cdots,x_{n})|\leq d_{i}$ (94)
for every $1\leq i\leq n$, where $d_{1},\cdots,d_{n}$ are some nonnegative real constants. Then, for every $r>0$,
$\mathbb{P}\big(U-\mathbb{E}U\geq r\big)\leq\exp\bigg(-\frac{2r^{2}}{\sum_{i=1}^{n}d_{i}^{2}}\bigg).$ (95)
@@
The following lemma is necessary for the application of McDiarmid’s Inequality for the codewords over the AWGN channel, the proof can be found in Appendix E.
@@
Let $\mathcal{C}_{1}\triangleq\left\{\bm{x}_{1},\cdots,\bm{x}_{k},\cdots,\bm{x}_{M}\right\}$ and $\mathcal{C}_{2}\triangleq\{\bm{x}_{1},\cdots,\bm{x}^{\prime}_{k},\cdots,\bm{x}_{M}\}$ be two codebooks of codewords with length $n$, which have only one difference on $k$-th codeword. The codewords are of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are all in the support set of $\bar{P}_{n}^{n}$. If $Q^{(n)}_{\bm{Z},\bm{C}_{1}}$ and $Q^{(n)}_{\bm{Z},\bm{C}_{2}}$ are the mixture Gaussian distributions induced by $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ through the channel $W_{2}$, respectively. The following bound holds for the difference between $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}_{1}}\|Q^{(n)}_{\bm{Z},0})$ and $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}_{2}}\|Q^{(n)}_{\bm{Z},0})$.
$\begin{split}\big|\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}_{1}}\|Q^{(n)}_{\bm{Z},0})&-\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}_{2}}\|Q^{(n)}_{\bm{Z},0})\big|\leq\frac{4n\Psi(n)}{M\sigma^{2}}.\end{split}$ (96)
@@
Consider the AWGN channel $W_{2}$ defined in (6), if $\{\bm{x}_{w},w=1,\cdots,M\}$ are $M$ codewords randomly sampled from $\bar{P}_{n}^{(n)}$ and $Q^{(n)}_{\bm{Z},\bm{C}}(\bm{Z})=\frac{1}{M}\sum_{w=1}^{M}W_{2}(\bm{Z}|\bm{x}_{w})$ is the induced mixture distribution of the codebook through the channel $W_{2}$, then for any $\lambda>0$, the following inequality holds,
$\mathbb{P}\bigg(\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})-\mathbb{E}\big(\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})\big)\geq\lambda\bigg)$
@@
The following lemma is almost a restatement of Lemma 4 in [22]. It provide a sufficient condition that we can expurgate part of the codewords so that the remaining codewords constitute a family of codebooks such that its induced distribution is close to the distribution of noise, at the same time, the average decoding error probability of the second type for each of the codebook is low.
@@
With AWGN channels $W_{1}$ and $W_{2}$ defined in (5) and (6), respectively, let the codewords $\bm{x}_{sw},s=1,\cdots,K,w=1,\cdots,M$ be i.i.d sampled from $\bar{P}_{n}^{(n)}$, and denote $\epsilon^{(1)}_{sw}$ and $\epsilon^{(2)}_{sw}$ as the decoding error probability of the first type and second type as (39) and (40). For any subset $\mathcal{I}\subset\langle M\rangle\times\langle K\rangle$, define a mixture distribution $Q_{\bm{Z},\bm{C}}^{\mathcal{I}}$ as
$Q_{\bm{Z},\bm{C}}^{\mathcal{I}}\triangleq\frac{1}{|\mathcal{I}|}\sum_{(s,w)\in\mathcal{I}}W^{n}_{2}(\bm{Z}|\bm{x}_{sw})$ (98)
For this mixture distribution and every positive $\lambda_{1},\lambda_{2},\lambda_{3}$, define events
$\begin{split}\mathcal{E}_{1}\triangleq\left\{\underset{s}{\max}\frac{1}{M}\sum_{w=1}^{M}\varepsilon_{sw}^{(2)}\leq\lambda_{1}\frac{1}{n^{c_{0}}}\mathbb{E}_{\bar{Q}^{(n)}_{\bm{Y},1}}\bigg(\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\bigg)\right\}\end{split}$ (99)
@@
$\begin{split}\mathcal{E}_{2}\triangleq&\left\{\forall\mathcal{I}\subset\langle K\rangle\times\langle M\rangle\,\ \,\text{with}\,\ \,|\mathcal{I}|=\lambda_{2}MK,\right.\\ \mathbb{D}(Q_{\bm{Z},\bm{C}}^{\mathcal{I}}\|Q^{(n)}_{\bm{Z},0})&\leq\log\big(1+\frac{2^{\gamma_{1}+1}}{\lambda_{2}MK}\big)+An\cdot ae^{-bn}\\ +&\left.\mathbb{D}(\bar{P}_{n}^{(n)}\|P^{(n)}_{n})+2\mathbb{D}(Q^{(n)}_{\bm{Z},1}\|Q^{(n)}_{\bm{Z},0})+\lambda_{3}\right\}.\end{split}$ (100)
@@
$\mathbb{P}(\mathcal{E}_{1})$ (102)
$=$ $\mathbb{P}\bigg(\underset{s}{\max}\frac{1}{M}\sum_{w=1}^{M}\varepsilon_{sw}^{(2)}\leq\lambda_{1}\frac{1}{n^{c_{0}}}\mathbb{E}_{\bar{Q}^{(n)}_{\bm{Y},1}}\bigg(\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\bigg)\bigg)$
$=$ $\bigg(1-\mathbb{P}\big(\frac{1}{M}\sum_{w=1}^{M}\varepsilon_{sw}^{(2)}\geq\lambda_{1}\frac{1}{n^{c_{0}}}\mathbb{E}_{\bar{Q}^{(n)}_{\bm{Y},1}}\bigg(\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\bigg)\bigg)^{K}$
@@
$\begin{split}&\mathbb{P}(\mathcal{E}_{1}\cap\mathcal{E}_{2})\\ =&1-\mathbb{P}(\mathcal{E}^{c}_{1}\cup\mathcal{E}^{c}_{2})\\ \geq&1-\mathbb{P}(\mathcal{E}^{c}_{1})-\mathbb{P}(\mathcal{E}^{c}_{2})\\ =&\mathbb{P}(\mathcal{E}_{1})+\mathbb{P}(\mathcal{E}_{1})-1\\ \geq&(1-\frac{1}{\lambda_{1}})^{K}-\exp(MK\mathbb{H}_{b}(\lambda_{2}))\times\exp\bigg(-\frac{\lambda_{2}\lambda_{3}^{2}MK\sigma^{4}}{8n^{2}\Psi^{2}(n)}\bigg)\end{split}$ (105)
@@
Remark 11.
Different from Lemma 4 in [22], here the event $\mathcal{E}_{1}$ only bounds the average error probabilities of the second type for all sub-codebooks. We have derived upper bound on the maximal decoding error probability of the first type in Lemma 4.
We have illustrated the random coding scheme, decoding method in Section III-A. The upper bound on the first type of decoding error in Lemma 4 and the average of the second type of decoding error has been analyzed in Lemma 6. We also bounded the expectation of the discriminative metric - $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})$. A legitimate covert code should be based on the previous lemma, the choice of various parameters both reliability and covertness will be illustrated in our main results as next theorem. We will provide characterizations on the size $M_{n}$ of each codebook for covert communication over AWGN channels with sufficient large $n$ through proper choice of the power level $\Psi(n)$. The theorem also characterizes the necessary number of key bits $K_{n}$, and it is shown that $K_{n}$ depends not only on the noise level $\sigma^{2}$ at the adversary’s channel and covert parameter $\delta$, but also depends on the hyper-parameter of truncation $\mu$.
@@
For covert communication over AWGN channels $W_{1}$ and $W_{2}$ in (5) and (6), a sufficient condition for the existence of an $(M,K,n,\epsilon,\sigma,\delta)$ code is listed as the following three requirements
@@
    Reliability (maximal error probability)
    $\epsilon_{1}+\frac{\lambda_{1}}{1-\lambda_{2}}\frac{1}{n^{c_{0}}}\mathbb{E}_{\bar{Q}^{(n)}_{\bm{Y},1}}\bigg(\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\bigg)\leq\epsilon,$ (107)
@@
    $\begin{split}\log&\big(1+\frac{2^{\gamma_{1}+1}}{\lambda_{2}M_{n}K_{n}}\big)+An\cdot ae^{-bn}\\ +&\mathbb{D}(\bar{P}_{n}^{(n)}\|P^{(n)}_{n})+(1+\iota^{\prime})\mathbb{D}(Q^{(n)}_{\bm{Z},1}\|Q^{(n)}_{\bm{Z},0})\\ +&\lambda_{3}\leq\delta.\end{split}$ (108)
@@
    Choose intermediate parameters $\lambda_{1},\lambda_{2},\lambda_{3}$ as (1): $\lambda_{1}=n^{c_{1}}$, (2) $\lambda_{2}=1-n^{-c_{2}}$, (3): $\lambda_{3}=n^{-c_{3}}$ such that $c_{1},c_{2},c_{3}$ all positive, meanwhile, they satisfy $0<c_{1}<c_{0}$, $c_{1}+c_{2}<c_{0}$ and $0<2c_{3}<c_{2}-1$.